OpenStudy (anonymous):

can you please help me solve for 'a': 2/3(5 - a) = 2

OpenStudy (anonymous):

2(5-a)=6 (muliplied both sides by 3) 5-a=3 (divided both sides by 2) a=2 (rearranged to make a the subject)

OpenStudy (anonymous):

@ stroodle i don't think 2 is the answer since idf you sub it in you don't get 2, you get 2/9

OpenStudy (anonymous):

If you sub in 2, you get 2=2, which is true.

OpenStudy (anonymous):

a = 28/6

OpenStudy (anonymous):

2/3(5-2) = 2/3x 3 =2/9 not 2

OpenStudy (anonymous):

In that case it should be written as 2/(3(5-a))

OpenStudy (anonymous):

oh i see you had interpreted the question differently from me